OpenStudy (anonymous):

A bag contains 1 blue, 2 white, and 2 red marbles. Match the columns. Drag the letter after a compound event in the left column and drop it before the matching probability in the right column. Term: P(red and red), with replacement P(white and white), with replacement P(blue and red), with replacement P(white and red), with replacement Definition: A) 20% B) 16% C) 8% D) 10%

OpenStudy (sriram):

red and red=B white and white=B blue and red=C white and red=B

OpenStudy (anonymous):

no

OpenStudy (anonymous):

For the probability of a red and a red, let's think. There are 2 reds out of 5 total marbles. So, the probability of drawing a red marble is 2/5. We get to put the marble we draw back, so the second time we draw the chance is still 2/5. So, the probability for both events happening together is 2/5*2/5=4/25=.16=16% The probability of a white and a white is the same as a red and a red because there are 2/5 for both types of marbles, hence this probability is also 16%. The probability of a blue and then a red. The probability of a blue is 1/5 and then since we put it back the probability of pulling a red is 2/5. The total probability is then 1/5*2/5=2/25=.08=8% The probability of a white and a red. Well the chances for a white are 2/5 then we put that white marble back in and draw a red, which is a 2/5 chance. So that probability is 2/5*2/5=4/25=.16=16% Hope that helps some.